# Text2SQL-Flow: A Robust SQL-Aware Data Augmentation Framework for Text-to-SQL

### v2

@@ IV-C Comparison of Retrieval Strategies
As shown in the lower part of Table III, fine-tuning the retrieval model on high-quality synthetic data from SQLFlow-part using our retrieval strategy significantly boosts performance. Both the high-quality augmented data and the masked alignment retrieval method contribute to this improvement. Specifically, our Question-SQL alignment method without masking achieves 59.3%, 81.0%, and 82.1% accuracy on BIRD-dev, Spider-dev, and Spider-test, respectively, outperforming the baseline that relies on Question-SQL similarity retrieval with the base (non-finetuned, no mask) model. This indicates that fine-tuning with augmented data enables the model to select more relevant examples, thereby enhancing the in-context learning capability of downstream closed-source models for accurate SQL generation. The masking strategy is another key contributor to this improvement. Ablation studies confirm its effectiveness: for non-finetuned methods, masking consistently improves retrieval performance (e.g., DAIL selection accuracy on BIRD-dev increases from 58.8% to 59.4%). Similarly, when we fine-tune on SQLFlow-part and use masking during training, performance improves to 61.0%, 81.2%, and 83.5% on the three benchmarks, demonstrating that masking remains essential even after fine-tuning. By masking literal values in the training data, the model learns to focus on the structural correspondence between questions and SQL queries while ignoring superficial lexical differences. This enables it to retrieve examples whose SQL patterns closely match the unseen ground-truth query based solely on the input question—in a single step. In contrast, methods like DAIL require generating an intermediate SQL query during retrieval, which incurs additional API costs and risks propagating errors from imperfect intermediate generations.
IV-D Comparison of Training Data for Retrieval Model
@@
As shown in Figure 5, the model trained solely on the BIRD dataset performs reasonably well on the BIRD-dev but suffers a sharp drop in performance on the Spider-dev and Spider-test. This indicates that models trained on a single dataset struggle to generalize effectively to other task settings.
@@
However, simply mixing data from multiple sources does not necessarily enhance generalization. For example, training on both Spider and BIRD data for the Spider benchmark even leads to degraded performance. In contrast, the model trained on the SQLFlow-part combined with Spider and BIRD achieves better results, effectively mitigating overfitting and substantially improving cross-benchmark performance. This indicates that the benefit does not stem merely from larger data volume, but from the diversity and quality of the data generated by Text2SQL-Flow, which more effectively boosts model generalization. A comparison with SynSQL-part further validates this conclusion. As a fully synthetic dataset, SynSQL exhibits a data distribution that diverges significantly from real-world applications. Although it aims for generalization, its effectiveness on specific tasks remains limited, demanding massive data and computation for only modest gains. For instance, SynSQL-part achieves only slight improvement on the Spider benchmark, barely exceeding the Spider+BIRD combination. In contrast, SQLFlow-part delivers markedly better results, highlighting its superior data quality, distribution alignment, and task relevance.
@@ IV-E Case Study of Few-shot Retrieval
V Conclusion
In this work, we introduced Text2SQL-Flow, a robust SQL-aware data augmentation framework that enables the scalable generation of high-quality, structurally diverse Text-to-SQL examples from limited seed data. Leveraging this framework, we constructed SQLFlow, a large-scale dataset comprising 89,544 annotated instances. Experimental results show that models fine-tuned on SQLFlow consistently outperform those trained on existing comparable datasets, highlighting the critical role of data quality and structural diversity in enhancing model generalization and compositional reasoning. Furthermore, our proposed masked alignment retrieval method improves the in-context learning performance of closed-source large language models by enabling structure-aware example selection. This work establishes a data-centric foundation for future Text-to-SQL research and demonstrates the importance of high-quality data in data-centric AI.

### v3

@@ IV-E Comparison of Retrieval Strategies
We compare our retrieval method with several baselines, where the embedding model is trained on SQLFlow-part, Spider-train, and BIRD-train, using Qwen3-Embedding-0.6B as the base model. As shown in the lower part of Table III, fine-tuning the retrieval model on high-quality synthetic data from SQLFlow-part using our retrieval strategy significantly boosts performance. Both the high-quality augmented data and the masked alignment retrieval method contribute to this improvement. Specifically, our Question-SQL alignment method without masking achieves 59.3%, 81.0%, and 82.1% accuracy on BIRD-dev, Spider-dev, and Spider-test, respectively, outperforming the baseline that relies on Question-SQL similarity retrieval with the base (non-finetuned, no mask) model. This indicates that fine-tuning with augmented data enables the model to select more relevant examples, thereby enhancing the in-context learning capability of downstream closed-source models for accurate SQL generation.
The masking strategy is another key contributor to this improvement. Ablation studies confirm its effectiveness: for non-finetuned methods, masking consistently improves retrieval performance (e.g., DAIL selection accuracy on BIRD-dev increases from 58.8% to 59.4%). Similarly, when we fine-tune on SQLFlow-part and use masking during training, performance improves to 61.0%, 81.2%, and 83.5% on the three benchmarks. By masking literal values in the training data, the model learns to focus on the structural correspondence between questions and SQL queries while ignoring superficial lexical differences. This enables it to retrieve examples whose SQL patterns closely match the unseen ground-truth query based solely on the input question in a single step. We also evaluate ASTs as an alternative SQL representation for embedding-model training and example retrieval in our Question-AST similarity setting. However, converting SQL into ASTs can obscure semantic cues but may introduce additional noise, leading to fewer improvements on some benchmarks. Since SQL already encodes rich structural information, our Question-SQL alignment is designed to enable the retrieval model to learn question–query correspondences in a data-driven manner while preserving the full SQL signal. In contrast, strategies such as DAIL Selection and PreSQL-SQL Selection require generating an intermediate SQL query for retrieval, which incurs additional API costs and risks propagating errors from intermediate generations.
IV-F Comparison of Training Data for Retrieval Model
@@
As shown in Figure 5, the model trained solely on the BIRD dataset performs well on BIRD-dev but suffers a sharp performance drop on Spider-dev and Spider-test, indicating limited cross-benchmark generalization when training on a single dataset. However, simply mixing data from multiple sources does not necessarily enhance generalization. For example, training on both Spider and BIRD data for the Spider benchmark even leads to degraded performance. In contrast, the model trained on the SQLFlow-part combined with Spider and BIRD achieves better results, effectively mitigating overfitting and substantially improving cross-benchmark performance. This indicates that the benefit does not stem merely from larger data volume, but from the diversity and quality of the data generated by Text2SQL-Flow, which more effectively boosts model generalization. A comparison with SynSQL-part further validates this conclusion. As a fully synthetic dataset, SynSQL exhibits a data distribution that diverges significantly from real-world applications. Although it aims for generalization, its effectiveness on specific tasks remains limited, demanding massive data and computation for only modest gains. For instance, SynSQL-part achieves only slight improvement on the Spider benchmark, barely exceeding the Spider+BIRD combination. In contrast, SQLFlow-part delivers markedly better results, highlighting its superior data quality, distribution alignment, and task relevance.
V Conclusion
In this work, we introduced Text2SQL-Flow, a robust SQL-aware data augmentation framework that enables the scalable generation of high-quality, structurally diverse Text-to-SQL examples from limited seed data. Leveraging this framework, we constructed SQLFlow, a large-scale dataset comprising 75,386 annotated instances. Experimental results show that models fine-tuned on SQLFlow consistently outperform those trained on existing comparable datasets, highlighting the critical role of data quality and structural diversity in enhancing model generalization and compositional reasoning. Furthermore, our proposed masked alignment retrieval method improves the in-context learning performance of closed-source large language models by enabling structure-aware example selection. This work establishes a data-centric foundation for future Text-to-SQL research and demonstrates the importance of high-quality data in data-centric AI.
Appendix A Additional Pipeline Components
In practical applications, to facilitate the organization and management of data generated by the pipeline and to support downstream usage, we additionally introduce two classifier steps to categorize the generated SQL queries. These operators are primarily user-oriented and are designed to improve system usability and engineering practicality, rather than to directly enhance the model performance emphasized in this paper. For this reason, their detailed descriptions are deferred to the Appendix. Based on the resulting classifications, users can more conveniently and efficiently inspect, analyze, and utilize the generated data. We describe the two operators as follows.
A-A SQL Component Difficulty Classifier
$cd_{i}=\text{CC}(s_{i}^{\text{aug}})$ (9)
where $\text{CC}(\cdot)$ denotes the mapping function of the classifier, and $cd_{i}$ represents the component difficulty of the query.
A-B SQL Execution Difficulty Classifier
$ed_{i}=1-\frac{n}{k}$ (10)
A-C Pipeline Workflow
In the complete pipeline, the two classifier operators are incorporated as integral components. For each generated SQL query, the SQL Component Classifier and the SQL Execution Classifier are applied to assign difficulty labels to the corresponding data sample. Through this process, the pipeline produces data samples in the following format:
Appendix B Details of Database Interaction Implementation
Within the framework of Text2SQL-Flow, a robust and efficient database interaction mechanism constitutes a core infrastructural component that underpins stable and scalable pipeline execution. Multiple critical stages of the pipeline, including SQL execution filtering and SQL augmentation, require frequent and reliable access to the underlying database system. These interactions can be broadly categorized into two types: (1) SQL Execution and (2) Schema Metadata Extraction.
SQL Execution refers to submitting generated SQL queries to the database engine and collecting execution feedback, which is primarily used to validate the executability of candidate queries. In practice, mainstream database systems such as SQLite, MySQL, and PostgreSQL expose heterogeneous driver interfaces and connection protocols. Directly relying on database-specific native APIs complicates cross-platform support and requires maintaining separate adaptation logic for each backend, leading to code redundancy and severely limiting system extensibility.
Schema Metadata Extraction aims to retrieve structured information about the database schema, including CREATE TABLE definitions, column attributes (e.g., data types, nullability, default values), primary and foreign key constraints, index metadata, and representative INSERT statements used for SQL augmentation. Such metadata is essential for understanding database semantics, generating syntactically and semantically valid SQL queries, and enabling context-aware data augmentation strategies. Although most relational databases expose schema information via mechanisms such as INFORMATION_SCHEMA or system catalog tables, their query syntax, naming conventions, and output formats vary substantially across systems. Moreover, for a given database instance, schema information typically remains stable over short time horizons. Repeatedly querying identical metadata therefore introduces unnecessary I/O overhead and latency, which can significantly degrade pipeline throughput at scale.
To address these challenges, we design and implement a Database Manager module as part of the data augmentation framework. This module abstracts low-level database interaction details and exposes a unified, efficient, and extensible programming interface to upper-layer components. Specifically, the Database Manager supports batch_sql_execution and batch_compare_sql, enabling higher throughput under high-concurrency workloads. It also encapsulates schema metadata retrieval and provides high-level interfaces such as get_ddl (for Data Definition Language extraction) and get_insert_statement (for generating representative insertion statements), thereby decoupling upper-layer logic from database-specific handling.
To ensure robust cross-database compatibility, we introduce an abstract base class, DatabaseConnector, which defines a standardized interaction interface. This interface includes methods such as connect_db (for connection establishment and initialization), execute_sql (for executing arbitrary SQL statements), and get_schema (for extracting complete schema metadata). Concrete connectors for specific database systems (e.g., SQLiteDatabaseConnector) inherit from this base class and implement database-specific driver calls and error handling. At runtime, the Database Manager dynamically instantiates the appropriate connector based on configuration, while its core logic remains agnostic to the underlying database type. This design achieves the architectural principle of “design once, adapt to many databases.”
To further improve throughput in large-scale SQL generation and validation scenarios, the Database Manager caches each database’s structured schema representation in memory. Although individual metadata queries are relatively fast, the pipeline repeatedly accesses the same schema when generating and validating multiple SQL candidates per database, making the cumulative overhead non-trivial. Upon the first access to a database, its schema (including tables, columns, keys, and sampled rows) is extracted, serialized, and stored in an in-memory cache. Subsequent accesses are served directly from the cache, avoiding repeated schema reconstruction and redundant metadata queries.
In summary, the Database Manager decouples the pipeline from underlying database implementations, optimizes execution efficiency, and enhances system modularity. Its extensible, plugin-style architecture lays a solid foundation for supporting additional database types in the future.
Appendix C Case Study of Few-shot Retrieval